# RASPNet: A Benchmark Dataset for Radar Adaptive Signal Processing Applications

### v2

@@ Abstract
We present a large-scale dataset for radar adaptive signal processing (RASP) applications to support the development of data-driven models within the adaptive radar community. The dataset, RASPNet, exceeds 16 TB in size and comprises 100 realistic scenarios compiled over a variety of topographies and land types from across the contiguous United States. For each scenario, RASPNet consists of 10,000 clutter realizations from an airborne radar setting, which can be used to benchmark radar and complex-valued learning algorithms. RASPNet intends to fill a prominent gap in the availability of a large-scale, realistic dataset that standardizes the evaluation of adaptive radar processing techniques and complex-valued neural networks. We outline its construction, organization, and several applications, including a transfer learning example to demonstrate how RASPNet can be used for realistic adaptive radar processing scenarios.
complex-valued data, deep learning, data-driven radar
@@ 1 Introduction
The success of radar technologies over the past century has been characterized by landmark advances in signal processing techniques. In this domain of radar adaptive signal processing (RASP), Space-Time Adaptive Processing (STAP) has been instrumental in achieving order-of-magnitude improvements in detection and localization performance, especially in environments dominated by interference and noise, which has revolutionized modern radar systems (Brennan & Reed, 1973). For decades, STAP has been integral to advancing surveillance, defense, and navigation (Ward, 1995; Li & Stoica, 2008). The foundations of RASP, realized using STAP, lies in the adaptive combination of spatial and temporal processing to optimize radar response to signals (Ward, 1998; Guerci, 2000; Melvin, 2004; Wicks et al., 2006).
@@
Despite the theoretical and practical advancements in RASP, the development, testing, and validation of RASP algorithms remains a significant challenge. This difficulty arises from the complexity of the real-world scenarios in which radar systems operate, which are characterized by dynamic and unpredictable environmental conditions. Theoretical models alone cannot capture the full range of operational scenarios, limiting the effectiveness of RASP algorithms developed in a vacuum (Klemm, 2002; Melvin, 2004).
@@
Paralleling the evolution of adaptive radar, complex-valued neural networks (CVNNs) have emerged as promising tools for enhancing signal processing capabilities (Hirose et al., 2006; Trabelsi et al., 2017; Wu et al., 2023). While CVNNs are well-suited for handling the inherently complex-valued data produced by radar systems, benchmarking CVNNs is limited by a lack of comprehensive datasets with naturally occurring complex-valued features (Mohammadi Asiyabi et al., 2023). Practitioners often convert real-valued datasets to the frequency domain for benchmarking (Trabelsi et al., 2017; Scardapane et al., 2020), which fails to capture the nuanced characteristics of real-world complex-valued data.
@@
The need for large-scale, comprehensive, realistic datasets in the adaptive radar community cannot be overstated. High-quality datasets serve multiple critical functions in the development lifecycle of RASP algorithms. Firstly, they provide a benchmark for evaluating and comparing the performance of algorithms under a wide range of conditions that mirror real-world operational environments. Furthermore, these datasets facilitate the development of data-driven approaches that leverage machine learning and artificial intelligence (Hara et al., 1994; Jiang et al., 2022; Venkatasubramanian et al., 2023). The Knowledge Aided Sensor Signal Processing and Expert Reasoning (KASSPER) dataset represents a seminal effort in this direction (Guerci & Baranoski, 2006), offering a focused repository of radar data that has been significant in advancing the field (Crooks, 2004; Blunt et al., 2006; Stoica et al., 2008; Wang et al., 2009).
@@
In spite of the value provided by existing repositories like KASSPER, the adaptive radar community continues to face data-related challenges. KASSPER, for instance, comprises one radar scenario, which is inadequate for data-driven algorithms. Thus, the primary issue is the absence of a publicly available dataset that encompasses the diversity of scenarios that adaptive radar systems encounter, as existing datasets are either too narrow in scope or restricted in accessibility and utility (Gogineni et al., 2022).
@@
To address these challenges, we present RASPNet, a large-scale, benchmark dataset for radar adaptive signal processing applications. RASPNet comprises a comprehensive set of $100$ scenarios from across the United States, compiled over a variety of environmental conditions, topographies, and land types via the RFView® modeling and simulation software. For each scenario, we provide $10{,}000$ realizations of clutter returns gathered with an airborne radar platform over a $20$ kilometer distance. RASPNet has the flexibility to support both traditional RASP algorithms and emerging data-driven approaches, facilitating a concrete benchmark to compare the performance of adaptive radar processing methods, and providing a critical testbed for CVNNs.
@@
The outline of this paper is as follows. In Section 2, we discuss the properties of RASPNet, including the radar parameters used to generate the dataset, the $100$ radar scenarios in the dataset, and validate RFView® versus measured data. In Section 3, we describe the organization of RASPNet via the energy distance metric. In Section 4, we delineate two relevant applications of RASPNet for benchmarking target localization accuracy, and for transfer learning in realistic adaptive radar processing scenarios. Finally, in Section 6, we summarize RASPNet and discuss future work. RASPNet is accessible at https://shyamven.github.io/RASPNet.
@@ 2.2 Radar Scenarios
Figure 1: Mapping the $M=100$ scenarios within RASPNet. Each scenario is indexed by $i\in S$, ordered westernmost to easternmost.
To evaluate whether RASPNet encompasses a realistic and extensive feature set, we empirically verify the exemplar nature of our $M$ hand-picked scenarios in Section 2.3. Additionally contained within these $M$ scenarios are $30$ land-to-water transition scenarios, which capture the complexities associated with implementing adaptive radar processing algorithms in coastal regions. The complete tabulation of scenarios comprising RASPNet, along with short descriptions, is provided in Section G of the Supplementary Information.
Algorithm 1 : Partitioning Around Medoids (PAM) with the $\mathcal{E}$-statistic
0:  $\,\overline{\!{S}}$ (set of indices), $M^{*}$ (number of representative indices), $T$ (max steps)
0:  $S^{*}$ (set of representative indices obtained through PAM)
1:  Initialize $E\leftarrow\mathbf{0}_{\,\overline{\!{M}}\times\,\overline{\!{M}}}$Initialize the matrix of $\mathcal{E}$-statistics}
2:  for $i=1$ to $\,\overline{\!{M}}$ do
3:     for $k=i$ to $\,\overline{\!{M}}$ do
4:        $E_{i,k}\leftarrow\mathcal{E}(Z^{i},Z^{k})$; $E_{k,i}\leftarrow\mathcal{E}(Z^{i},Z^{k})$Compute and store the $\mathcal{E}$-statistic for the pair $(i,k)$}
5:     end for
6:  end for
7:  $S^{*}\leftarrow\text{SRSWOR}(\,\overline{\!{S}},M^{*})$; $C\leftarrow\sum_{i\in\,\overline{\!{S}}}\min_{k\in S^{*}}E_{i,k}$Initialize $S^{*}$ and compute initial cost $C$}
8:  for $t=1$ to $T$ do
9:     for $m\in S^{*}$ do
10:        for $o\in\,\overline{\!{S}}\setminus S^{*}$ do
11:           $S^{\prime*}\leftarrow(S^{*}\setminus\{m\})\cup\{o\}$; $C^{\prime}\leftarrow\sum_{i\in\,\overline{\!{S}}}\min_{k\in S^{\prime*}}E_{i,k}$Form $S^{\prime*}$ and compute new cost $C^{\prime}$}
12:           if $C^{\prime}<C$ then
13:              $S^{*}\leftarrow S^{\prime*}$; $C\leftarrow C^{\prime}$Update $S^{*}$ and $C$ to the new values}
14:           end if
15:        end for
16:     end for
17:  end for
18:  return $S^{*}$
@@ 2.3 Exemplar Scenario Validation
To determine whether our $100$ hand-picked scenarios are exemplar, we compare their geographic coverage with that of a collection of $100$ scenarios obtained through the Partitioning Around Medoids (PAM) algorithm (Kaufman & Rousseeuw, 1990), from $500$ total scenarios uniformly selected at random across the contiguous United States. We denote $\smash{\,\overline{\!{S}}=\{1,2,\ldots,\,\overline{\!{M}}\}}$ as the set of indices pertaining to the $\smash{\,\overline{\!{M}}=500}$ scenarios, and denote $\smash{S^{*}\subset\,\overline{\!{S}}}$ as the set of indices yielded by PAM, where $\left|S^{*}\right|=M^{*}=100$.
To determine $S^{*}$ using the PAM algorithm, we employ the energy statistic ($\mathcal{E}$-statistic) as our distance metric (Székely & Rizzo, 2013). The $\mathcal{E}$-statistic quantifies the statistical distance between the distributions of random variables by aggregating pairwise sample distances, capturing differences in their distributional characteristics. Rooted in potential energy principles, the $\mathcal{E}$-statistic facilitates robust comparisons of complex, multivariate datasets beyond location or scale. The formal definition of the $\mathcal{E}$-statistic for our analysis is provided in Definition 2.1.
@@
$\mathcal{E}(Z^{i},Z^{k})=\frac{2}{K^{2}}\sum_{j=1}^{K}\sum_{l=1}^{K}\|z^{i}_{j}-z^{k}_{l}\|$ (1)
$-\frac{1}{K^{2}}\sum_{j=1}^{K}\sum_{l=1}^{K}\|z^{i}_{j}-z^{i}_{l}\|-\frac{1}{K^{2}}\sum_{j=1}^{K}\sum_{l=1}^{K}\|z^{k}_{j}-z^{k}_{l}\|.$
[…]
We observe that the geographic coverage of the $100$ hand-picked scenarios in RASPNet closely mirrors the coverage of the $100$ representative scenarios determined through PAM. A large concentration of these scenarios are in the Western United States — this region is characterized by a wider variety of terrain types, which introduces greater variability in the distribution of geographic features, leading to a greater concentration of representative scenarios. We also note that the $100$ hand-picked scenarios in RASPNet are more dispersed across the contiguous United States by construction, particularly because they encompass identifiable geographical landmarks representing a broad spectrum of environmental conditions. A greater emphasis was also placed on incorporating land-to-water transition zones in the $100$ hand-picked scenarios, acknowledging these as inherently more varied scenarios.
@@
Figure 2: Range-Doppler plots for (a) Flight #5 Acquisition 575 from the MCARM measured dataset, (b) RFView® simulated data, and (c) classical bald-earth model without terrain heights. The colormap depicts the magnitude squared of the radar range-Doppler spectrum, with red corresponding to higher magnitudes and blue to lower magnitudes.
2.4 Validation Against Measured Data
[…]
To validate the similarity between the RFView® simulated data and the MCARM measured data, we perform a modified version of the likelihood-based test from (Johnson & Abramovich, 2010). Let $\smash{\mathbf{Z}^{\text{MCM}}[l,g]\in\mathbb{C}^{\tilde{L}\times 1}}$, $\smash{\mathbf{Z}^{\text{RF}}[l,g]}$, and $\smash{\mathbf{Z}^{\text{BE}}[l,g]\in\mathbb{C}^{\tilde{L}\times K}}$ form the MCARM, RFView®, and bald-earth clutter realizations, for the $l$-th Doppler bin and $g$-th range bin. The null hypothesis, $H_{0}$, asserts that $\mathbf{Z}^{\text{MCM}}[l,g]$ is from the same distribution as $\mathbf{Z}^{\text{RF}}[l,g]$ — we repeat this procedure for $\mathbf{Z}^{\text{BE}}[l,g]$). Next, $\forall l,g$, we compare the likelihood of the MCARM clutter realization to the empirical distribution of likelihoods from the RFView® clutter realizations. We let $\mathbf{\Sigma}_{\text{RF}}[l,g]$ denote the sample covariance matrix for the RFView case, where the likelihood of the MCARM data given $\mathbf{\Sigma}_{\text{RF}}[l,g]$, and the likelihood of the $j^{\text{th}}$ RFView clutter realization given $\mathbf{\Sigma}_{\text{RF}}[l,g]$, are given by:
$\mathcal{L}_{\text{MCM}}[l,g]$ $=\frac{e^{\left(-\frac{1}{2}\mathbf{Z}^{\text{MCM}}[l,g]^{H}\boldsymbol{\Sigma}_{\text{RF}}[l,g]^{-1}\mathbf{Z}^{\text{MCM}}[l,g]\right)}}{\sqrt{(2\pi)^{L}\det(\boldsymbol{\Sigma}_{\text{RF}}[l,g])}},$ (2)
$\mathcal{L}_{\text{RF},j}[l,g]$ $=\frac{e^{\left(-\frac{1}{2}\mathbf{Z}_{j}^{\text{RF}}[l,g]^{H}\boldsymbol{\Sigma}_{\text{RF}}[l,g]^{-1}\mathbf{Z}_{j}^{\text{RF}}[l,g]\right)}}{\sqrt{(2\pi)^{L}\det(\boldsymbol{\Sigma}_{\text{RF}}[l,g])}}.$ (3)
@@ 2.4 Validation Against Measured Data
$\rho=\frac{1}{\Lambda G}\sum_{l=1}^{\Lambda}\sum_{g=1}^{G}\mathbb{I}\big{(}\mathcal{L}_{\text{MCM}}[l,g]\in$ (4)
$\quad\quad\quad\big{[}\text{\textperthousand}_{2.5}(\mathcal{L}_{\text{RF}}[l,g]),\text{\textperthousand}_{97.5}(\mathcal{L}_{\text{RF}}[l,g])\big{]}\big{)}.$

### v3

@@ Abstract
We present a large-scale dataset called RASPNet for radar adaptive signal processing (RASP) applications to support the development of data-driven models within the adaptive radar community. RASPNet exceeds 16 TB in size and comprises 100 realistic scenarios compiled over a variety of topographies and land types across the contiguous United States. For each scenario, RASPNet comprises 10,000 clutter realizations from an airborne radar setting, which can be used to benchmark radar and complex-valued learning algorithms. RASPNet intends to fill a prominent gap in the availability of a large-scale, realistic dataset that standardizes the evaluation of RASP techniques and complex-valued neural networks. We outline its construction, organization, and several applications, including a transfer learning example to demonstrate how RASPNet can be used for real-world adaptive radar scenarios.
footnotetext: GitHub Repository: https://github.com/shyamven/RASPNet.
@@ 1 Introduction
The success of radar technologies over the past century has been characterized by landmark advances in radar adaptive signal processing (RASP). Within RASP, Space-Time Adaptive Processing (STAP) has been instrumental in achieving order-of-magnitude improvements in detection and localization performance, especially in environments dominated by clutter and noise, which has revolutionized modern radar systems (Brennan and Reed, 1973). For decades, STAP has been integral to advancing surveillance, defense, and navigation (Ward, 1995; Li and Stoica, 2008). The foundations of RASP, realized using STAP, lies in the adaptive combination of spatial and temporal processing to optimize radar response to signals (Ward, 1998; Guerci, 2000; Melvin, 2004; Wicks et al., 2006).
@@
Despite the theoretical advancements in RASP, the development, testing, and validation of RASP algorithms remains a significant challenge. This difficulty arises from the complexity of the real-world scenarios in which radar systems operate, which are characterized by dynamic and unpredictable environmental conditions. Theoretical models cannot capture the full range of operational scenarios, limiting the effectiveness of RASP algorithms developed in a vacuum (Klemm, 2002; Melvin, 2004).
@@
Paralleling the evolution of adaptive radar, complex-valued neural networks (CVNNs) have emerged as promising tools for enhancing signal processing capabilities (Hirose et al., 2006; Trabelsi et al., 2017; Wu et al., 2023). While CVNNs are well-suited for handling the inherently complex-valued data produced by radar systems, benchmarking CVNNs is limited by a lack of comprehensive datasets with naturally occurring complex-valued features (Mohammadi Asiyabi et al., 2023). Practitioners often convert real-valued datasets to the frequency domain (Trabelsi et al., 2017; Scardapane et al., 2020), which fails to capture the nuanced characteristics of real-world complex-valued data.
@@
The need for large-scale, comprehensive, realistic datasets in the adaptive radar community cannot be overstated. High-quality datasets serve critical functions in the development of RASP algorithms. Firstly, they provide a benchmark for evaluating and comparing the performance of algorithms under conditions that mirror real-world operational environments. Furthermore, these datasets facilitate the development of data-driven approaches that leverage machine learning (Hara et al., 1994; Jiang et al., 2022; Venkatasubramanian et al., 2024). The Knowledge Aided Sensor Signal Processing and Expert Reasoning (KASSPER) dataset represents a seminal effort in this direction (Guerci and Baranoski, 2006), offering a focused repository of radar data that has been significant in advancing the field (Crooks, 2004; Blunt et al., 2006; Stoica et al., 2008; Wang et al., 2009).
@@
In spite of the value provided by existing repositories like KASSPER, the adaptive radar community faces significant data-related challenges. KASSPER, for instance, comprises one 5 MB radar scenario, which is inadequate for data-driven algorithms. Thus, the primary issue is the absence of a publicly available dataset that encompasses the diversity of scenarios that adaptive radar systems encounter, as existing datasets are either too narrow in scope or have restricted access (Gogineni et al., 2022).
@@
To address these challenges, we present RASPNet, a large-scale, benchmark dataset for radar adaptive signal processing applications. Eclipsing $16$ TB in size, RASPNet comprises a comprehensive set of $100$ radar scenarios from across the United States, compiled over various environmental conditions and land types via the RFView® modeling software. For each scenario, we provide 10,00010{,}000 realizations of clutter returns gathered with an airborne radar platform. RASPNet supports both traditional RASP algorithms and emerging data-driven approaches, facilitating a concrete benchmark to compare the performance of adaptive radar processing methods, and providing a critical testbed for CVNNs.
@@
The outline of this paper is as follows. In Section 2, we discuss the properties of RASPNet, including the radar parameters used to generate the dataset, the $100$ radar scenarios in the dataset, and validate RFView® versus measured data. In Section 3, we describe the organization of RASPNet via the energy distance metric. In Section 4, we delineate several applications of RASPNet for benchmarking target localization accuracy, benchmarking complex-valued learning algorithms, and for transfer learning in realistic adaptive radar processing scenarios. In particular, we propose a convolutional transformer architecture, termed the Adaptive Radar Transformer, which achieves benchmark target localization accuracy on RASPNet, and we extend the Steinmetz network construction from (Venkatasubramanian et al., 2025) to propose the Convolutional Steinmetz Network, which achieves benchmark performance on the RASPNet CVNN task. Finally, in Section 6, we summarize RASPNet and discuss future work. RASPNet is accessible at: https://www.sdms.afrl.af.mil/index.php?collection=raspnet.
@@ 2.2 Radar Scenarios
Figure 1: Mapping the $M=100$ scenarios within RASPNet. Each scenario is indexed by $i\in S$.
To evaluate the coverage of RASPNet scenarios, we empirically verify the exemplar nature of our $M$ hand-picked scenarios in Section 2.3. The complete tabulation of scenarios comprising RASPNet, along with short descriptions, is provided in Section G of the Supplementary Information.
@@ 2.3 Exemplar Scenario Validation
To determine whether our $100$ hand-picked scenarios are exemplar, we compare their geographic coverage with that of a collection of $100$ scenarios obtained through the Partitioning Around Medoids (PAM) algorithm (Kaufman and Rousseeuw, 1990), from $500$ total scenarios uniformly selected at random across the United States. We denote $\smash{\,\overline{\!{S}}=\{1,2,\ldots,\,\overline{\!{M}}\}}$ as the set of indices pertaining to the $\smash{\,\overline{\!{M}}=500}$ scenarios, and $\smash{S^{*}\subset\,\overline{\!{S}}}$ as the set of indices yielded by PAM, where $\left|S^{*}\right|=M^{*}=100$.
Input: $\overline{\!{S}}$ (set of indices), $M^{*}$ (number of representative indices), $T$ (max steps)
Output: $S^{*}$ (set of representative indices obtained through PAM)
1
[…]
3 for $i=1$ to $\overline{\!{M}}$ do
4 for $k=i$ to $\overline{\!{M}}$ do $E_{i,k}\leftarrow\mathcal{E}(Z^{i},Z^{k})$; $E_{k,i}\leftarrow\mathcal{E}(Z^{i},Z^{k})$ ;
5
6 end
[…]
8 for $t=1$ to $T$ do
9 for $m\in S^{*}$ do
10    for $o\in\,\overline{\!{S}}\setminus S^{*}$ do
11       $S^{\prime*}\leftarrow(S^{*}\setminus\{m\})\cup\{o\}$; $C^{\prime}\leftarrow\sum_{i\in\,\overline{\!{S}}}\min_{k\in S^{\prime*}}E_{i,k}$;
12       if $C^{\prime}<C$ then $S^{*}\leftarrow S^{\prime*}$; $C\leftarrow C^{\prime}$ ;
13
14      end
15
16   end
17
18 end
return $S^{*}$;
Algorithm 1 Partitioning Around Medoids (PAM) with the $\mathcal{E}$-statistic
To determine $S^{*}$ using the PAM algorithm, we employ the energy statistic ($\mathcal{E}$-statistic) as our distance metric (Székely and Rizzo, 2013). The $\mathcal{E}$-statistic characterizes the statistical distance between the distributions of random variables by aggregating pairwise sample distances, capturing differences in their distributional characteristics. Rooted in potential energy principles, the $\mathcal{E}$-statistic facilitates robust comparisons of complex, multivariate datasets beyond location or scale. The formal definition of the $\mathcal{E}$-statistic for our analysis is provided in Definition 2.1.
@@
$\mathcal{E}(Z^{i},Z^{k})=\frac{2}{K^{2}}\sum_{j=1}^{K}\sum_{l=1}^{K}\|z^{i}_{j}-z^{k}_{l}\|-\frac{1}{K^{2}}\sum_{j=1}^{K}\sum_{l=1}^{K}\|z^{i}_{j}-z^{i}_{l}\|-\frac{1}{K^{2}}\sum_{j=1}^{K}\sum_{l=1}^{K}\|z^{k}_{j}-z^{k}_{l}\|.$ (1)
We now outline the PAM algorithm for our analysis. We first construct the $\smash{\,\overline{\!{M}}\times\,\overline{\!{M}}}$ $\mathcal{E}$-statistic matrix, $E$, for $\smash{\,\overline{\!{M}}=500}$, where each element, $E_{i,k}$, is precomputed using $\mathcal{E}(Z^{i},Z^{k})$ for every pair $\smash{i,k\in\,\overline{\!{S}}}$. This precomputation step facilitates efficient distance lookups during the PAM algorithm’s execution. Over $T$ steps, the PAM algorithm iteratively refines a subset of $M^{*}=100$ representative scenarios, $S^{*}$, through a process of swapping current medoids with non-medoids to minimize the total cost, $C$, where $C$ is computed as the sum of distances from each scenario to its nearest medoid, based on the precomputed $\mathcal{E}$-statistic matrix, $E$. This procedure is summarized in Algorithm 1. The $500$ total scenarios and the $100$ representative scenarios from PAM are depicted in Section A of the Appendix.
We observe that the geographic coverage of the $100$ hand-picked scenarios in RASPNet closely mirrors the coverage of the $100$ representative scenarios determined through PAM. A large concentration of these scenarios are in the Western United States — this region is characterized by a wider variety of terrain types, which introduces greater variability in the distribution of geographic features, leading to a greater concentration of representative scenarios. We also note that the $100$ hand-picked scenarios in RASPNet are more dispersed across the contiguous United States by construction, as they encompass identifiable geographical landmarks representing a broad spectrum of environmental conditions.
2.4 Validation Against Measured Data
To evaluate the realistic nature of the RFView® simulator, we compare it with the MCARM measured dataset, focusing on Flight #5 Acquisition 575 (Babu et al., 1996). The MCARM radar parameters are tabulated in Section B of the Appendix. Let $\smash{Z^{i}\in\mathbb{C}^{\tilde{\Lambda}\tilde{L}\tilde{G}}}$, for $\smash{i\in\hat{S}=\{\text{MCM, RF, BE}\}}$, denote the clutter random variables for the MCARM, RFView®, and bald-earth (BE) models, wherein the bald-earth model is a more classical clutter model without terrain heights. For the RFView® and BE models, we generate the clutter datasets $\smash{\{z^{\text{RF}}_{j},\,\forall j\in D\}}$ and $\smash{\{z^{\text{BE}}_{j},\,\forall j\in D\}}$, which we then reshape to obtain $\smash{\mathbf{Z}^{\text{RF}},\mathbf{Z}^{\text{BE}}\in\mathbb{C}^{\tilde{\Lambda}\tilde{L}\tilde{G}\times K}}$. The MCARM Flight #5 Acquisition 575 provides a single clutter realization, $\smash{z^{\text{MCM}}_{1}}$, which we reshape to form $\smash{\mathbf{Z}^{\text{MCM}}\in\mathbb{C}^{\tilde{\Lambda}\tilde{L}\tilde{G}\times 1}}$.
@@ 2.4 Validation Against Measured Data
Figure 2: Range-Doppler plots for (a) Flight #5 Acquisition 575 from the MCARM measured dataset, (b) RFView® simulated data, and (c) classical bald-earth model without terrain heights. The colormap depicts the magnitude squared of the radar range-Doppler spectrum (red is higher and blue is lower).
To validate the similarity between the RFView® simulated data and the MCARM measured data, we perform a modified version of the likelihood-based test from (Johnson and Abramovich, 2010). Let $\smash{\mathbf{Z}^{\text{MCM}}[l,g]\in\mathbb{C}^{\tilde{L}\times 1}}$, $\smash{\mathbf{Z}^{\text{RF}}[l,g]}$, and $\smash{\mathbf{Z}^{\text{BE}}[l,g]\in\mathbb{C}^{\tilde{L}\times K}}$ form the MCARM, RFView®, and bald-earth clutter realizations, for the $l$-th Doppler bin and $g$-th range bin. The null hypothesis, $H_{0}$, asserts that $\mathbf{Z}^{\text{MCM}}[l,g]$ is from the same distribution as $\mathbf{Z}^{\text{RF}}[l,g]$ — we repeat this procedure for $\mathbf{Z}^{\text{BE}}[l,g]$). Next, $\forall l,g$, we compare the likelihood of the MCARM clutter realization to the empirical distribution of likelihoods from the RFView® clutter realizations. We let $\mathbf{\Sigma}_{\text{RF}}[l,g]$ denote the sample covariance matrix for the RFView case, such that the likelihood of the MCARM data given $\mathbf{\Sigma}_{\text{RF}}[l,g]$, and the likelihood of the $j^{\text{th}}$ RFView clutter realization given $\mathbf{\Sigma}_{\text{RF}}[l,g]$, are given by:
$\mathcal{L}_{\text{MCM}}[l,g]=\frac{e^{\left(-\frac{1}{2}\mathbf{Z}^{\text{MCM}}[l,g]^{H}\boldsymbol{\Sigma}_{\text{RF}}[l,g]^{-1}\mathbf{Z}^{\text{MCM}}[l,g]\right)}}{\sqrt{(2\pi)^{L}\det(\boldsymbol{\Sigma}_{\text{RF}}[l,g])}},\ \mathcal{L}_{\text{RF},j}[l,g]=\frac{e^{\left(-\frac{1}{2}\mathbf{Z}_{j}^{\text{RF}}[l,g]^{H}\boldsymbol{\Sigma}_{\text{RF}}[l,g]^{-1}\mathbf{Z}_{j}^{\text{RF}}[l,g]\right)}}{\sqrt{(2\pi)^{L}\det(\boldsymbol{\Sigma}_{\text{RF}}[l,g])}}.$ (2)
@@
$\rho=\frac{1}{\Lambda G}\sum_{l=1}^{\Lambda}\sum_{g=1}^{G}\mathbb{I}\big(\mathcal{L}_{\text{MCM}}[l,g]\in\big[\text{\textperthousand}_{2.5}(\mathcal{L}_{\text{RF}}[l,g]),\text{\textperthousand}_{97.5}(\mathcal{L}_{\text{RF}}[l,g])\big]\big).$ (3)